OpenStudy (anonymous):

Find the complete solution of the linear system or that it is inconsistent. -x+8y+5z=6 x -2z=0 4x-8y-11z=3

OpenStudy (anonymous):

if x-2z = 0 then x = 2z. Now you can use x = 2z, in -x+8y+5z = 6 and 4x-8y-11z = 3. You will get two equations with only y,z as the variables. Hence solve for y,z. Hence use x = 2z.